OpenStudy (anonymous):

find the sum of the interior angles and the number of diagonals of a regular polygon whose central angle is 6

Directrix (directrix):

If the regular polygon has a central angle measure of 6, then the number of sides of the polygon is 360/6 = 60. So, use the interior angle sum formula for polygons to get the sum for a 60-gon. Sum of Interior Angles = (n-2) × 180° where n is the number of sides of the polygon. @miggy9

OpenStudy (anonymous):

Thanks.

OpenStudy (jmark):

Given regular polygon whose central angle is 6 =>we have central angle =360/n =>n=360/6=60 number of diagonals =60 then sum of the interior angles =180(n-2) =>180(60-2) =>sum of the interior angles = 10440 Learn more about Irregular Polygon http://goo.gl/i2ZDYO